# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ VI-D Medium/Long Block-Lengths
(a) Sum-BER with medium/long block-lengths $L_{1},L_{2}$.
(b) Sum-BLER with medium block-length $L_{1}=L_{2}=120$.
Figure 8: Sum-error with medium/long block-lengths $L_{1},L_{2}$ where $\text{SNR}^{\text{ch}}_{1}=-1$dB and coding rate $r_{1}=r_{2}=1/3$. Under the high noise scenario with $\text{SNR}^{\text{ch}}_{1}=-1$dB, RNN-based coding performs well in terms of balancing the communication reliability when the channel noises become more asymmetric.
We next analyze a medium and long block-length regime, where we consider turbo coding [31] adopted in LTE standards as another baseline for medium/long block-length codes. We consider the trellis with $(4,[13,15])$, BPSK modulation, and $10$ decoding iterations. Fig. 8(7(a)) shows sum-BER curves under varying $\text{SNR}^{\text{ch}}_{2}$ where $\text{SNR}^{\text{ch}}_{1}=-1$dB and $r_{1}=r_{2}=1/3$. Since turbo coding exploits the block-length gain, a longer block of bits leads to a better sum-BER performance. However, for given $L_{1}$ and $L_{2}$, the sum-BER performances are almost constant over $\text{SNR}^{\text{ch}}_{2}$ since sum-BER is dominated by a fixed value of $\text{BER}_{1}$. On the other hand, our proposed two-way coding schemes, i.e., linear and RNN-based coding, significantly improve $\text{BER}_{1}$ as $\text{SNR}^{\text{ch}}_{2}$ increases, due to the two-way interactions between the users, which leads to substantial improvements in the sum-BER.
Fig. 8(7(b)) shows the sum-BLER performance for medium block-length $L_{1}=L_{2}=120$, where $\text{SNR}^{\text{ch}}_{1}=-1$dB and rate $r_{1}=r_{2}=1/3$. Since the SNR condition is poor, i.e., $\text{SNR}^{\text{ch}}_{1}=-1$dB, turbo coding does not perform well, resulting in a performance degradation in terms of sum-BLER. However, the two-way coding schemes, in particular RNN-based coding, improve the error performance due to their robustness to channel noises, and ultimately balance the communication reliability, i.e., obtain much smaller sum-BLER values.
@@ VI-E Threshold for Two-Way Coding Benefits
Figure 9: Each of BLER curves with block-length $L_{1}=L_{2}=3$ obtained by RNN-based coding with the sigmoid function and $K_{1}=K_{2}=3$ bits. The users start to benefit from two-way coding when $\text{SNR}^{\text{ch}}_{2}$ is larger than a threshold, which is 6dB in (a) and 5dB in (b), respectively.
@@
We have so far observed that our two-way coding schemes significantly improve the sum-error performance when the channel conditions are asymmetric. In this subsection, we further investigate each error curve to analyze how communication reliability can be balanced by our two-way coding scheme. Fig. 9 shows the BLER curves obtained by our RNN-based coding scheme with the sigmoid function and $K_{1}=K_{2}=3$ bits, where BLER is measured over the block-length $L_{1}=L_{2}=3$ bits. We consider $\text{SNR}^{\text{ch}}_{1}=1$dB in Fig. 9(8(a)) and $\text{SNR}^{\text{ch}}_{1}=-1$dB in Fig. 9(8(b)). To understand the interrelation between the two users through the proposed two-way coding, we introduce a baseline by modifying our RNN-based coding to generate two independent open-loop codes. Specifically, in Fig. 4, we provide $x_{i}[k-1]$ as an input to the encoder of User $i$ rather than $y_{i}[k-1]$ to remove the dependency between the encoders of the users, which is denoted by “RNN, open-loop” in the plots. Under the asymmetric channels with $\text{SNR}^{\text{ch}}_{1}<\text{SNR}^{\text{ch}}_{2}$, $\text{BLER}_{1}$ and $\text{BLER}_{2}$ obtained by “RNN, open-loop” is the upper and lower bound, respectively, of those obtained by our two-way coding scheme.
@@
In the figures, $\text{BLER}_{1}$ with open-loop coding yields a constant value, since $\text{BLER}_{1}$ is only determined by the fixed value of $\text{SNR}^{\text{ch}}_{1}$, while $\text{BLER}_{2}$ with open-loop coding decreases along $\text{SNR}^{\text{ch}}_{2}$. The results show that the two independent open-loop codes do not balance the communication reliability when the channels are asymmetric. However, our two-way coding scheme balances the communication reliability under asymmetric channels due to the interactive exchange of information between the users. In Fig. 9(8(a)), when $\text{SNR}^{\text{ch}}_{2}\geq 6$dB, $\text{BLER}_{1}$ is improved relative to that of the open-loop codes, while $\text{BLER}_{2}$ is sacrificed as compared to the open-loop coding. This is because User 2 (with higher SNR) helps User 1 (with lower SNR) through the interactive exchange of information to improve $\text{BLER}_{1}$ by implicitly providing feedback information to User 1. This behavior is in line with the insight discussed in [6] that a user with lower channel noise can act as a helper. Furthermore, we can identify a threshold of $\text{SNR}^{\text{ch}}_{2}$, above which the users start to benefit from the two-way interaction of our coding scheme. When $\text{SNR}^{\text{ch}}_{1}=1$dB, the threshold is $\text{SNR}^{\text{ch}}_{2}=6$dB as indicated in Fig. 9(8(a)), whereas, the threshold is $\text{SNR}^{\text{ch}}_{2}=5$dB in Fig. 9(8(b)) when $\text{SNR}^{\text{ch}}_{1}=-1$dB.
@@ VI-F The Superior Scheme over Different SNR Regions
Figure 10: The superior scheme in sum-BLER with block-lengths $L_{1}=L_{2}=6$ under various SNR scenarios. When SNRs are high and asymmetric, linear coding performs best, while RNN-based coding excels in low SNRs.
@@
Next, we explore different coding scheme among repetitive coding, TBCC, linear coding, and RNN-based coding in terms of sum-BLER for block-lengths $L_{1}=L_{2}=6$ under various SNR scenarios. The results are demonstrated in Fig. 10. In high and asymmetric SNR regions, we observe that linear coding outperforms the other schemes. On the other hand, in low SNR regions, RNN-based coding with the softmax function and $K_{1}=K_{2}=6$ bits yields the best performance, demonstrating its robustness to channel noises. Specifically, under symmetric SNR scenarios ($\text{SNR}^{\text{ch}}_{1}=\text{SNR}^{\text{ch}}_{2}=-5,-1,1,5$dB), RNN-based coding behaves like two independent open-loop coding when the channels are symmetric, as shown in Fig. 9, which outperforms TBCC. This result suggests that our learning architecture can be also utilized to develop error correction codes. While our study focuses on two-way coding behavior, the construction of error correcting codes is a potential area for future research. Note that, for moderate and asymmetric SNRs, RNN-based coding with the sigmoid function performs the best among the schemes.
VI-G Varying Coding Rates
@@
Figure 11: Sum-BER with various coding rates of $r_{2}$, where $\text{SNR}^{\text{ch}}_{1}=1$dB, $\text{SNR}^{\text{ch}}_{2}=20$dB, and $r_{1}=1/3$. RNN-based coding outperforms the counterparts in sum-BER even while supporting high coding rates.
@@
We so far looked into the benefit of two-way coding in terms of improving sum-error performance when coding rates are fixed. However, when $\text{SNR}^{\text{ch}}_{2}$ is high enough, while helping User 1 to improve sum-error, User 2 can increase its coding rate. To investigate the relationship between the coding rates and sum-error, we show the sum-BER performances with varying coding rates of $r_{2}$ in Fig. 11, where $r_{1}=1/3$, $\text{SNR}^{\text{ch}}_{1}=1$dB, and $\text{SNR}^{\text{ch}}_{2}=20$dB. For linear coding, we consider three different modulation bits at User 1, $K_{1}=1,2,3$. Correspondingly, we set the modulation bits at User 2 as $K_{2}=3r_{2},6r_{2},9r_{2}$. We consider the strategy of alternate channel uses discussed in Sec. VI-C for linear coding, where two message pairs are exchanged in an alternate manner. Specifically, $2K_{1}$ and $3K_{1}r_{2}$ bits are modulated to two messages at User $1$ denoted by $m_{1},m_{1}^{\prime}$ and at User $2$ denoted by $m_{2},m_{2}^{\prime}$, respectively, and the two pairs of messages, $(m_{1},m_{2})$ and $(m_{1}^{\prime},m_{2}^{\prime})$, are exchanged over $6K_{1}$ channel uses in an alternate manner. For RNN-based coding, the number of bits input to the encoders of Users 1 and 2 are $K_{1}=3$ and $K_{2}=9r_{2}$, respectively. From the figure, we observe that RNN-based coding outperforms the counterparts even under very high coding rates, $r_{2}\leq 3$, while linear coding outperforms the repetitive coding when the coding rate is quite low, $r_{2}\leq 1.5$.
@@ VI-H Power Distribution in Linear and RNN-Based Coding
We next investigate the power distribution from linear coding when adopting the strategy of alternate channel uses, where two users convey $L_{1}=L_{2}=6$ bits to one another over $N_{\text{L}}=18$ channel uses, where $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB. We consider two different modulation bits, $K_{1}=K_{2}=1$ in Fig. 12(11(a)) and $K_{1}=K_{2}=3$ in Fig. 12(11(b)). First, in Fig. 12(11(a)), each User $i$ has $M_{i}=L_{i}/K_{i}=6$ message symbols, where every two message pairs are transmitted over 6 channel uses through the alternate channel use strategy, as shown in Fig. 7(6(b)). Specifically, for the first message pair, the 1st, 3rd and 5th channel uses at User 1 and the 2nd, 4th and 6th channel uses at User 2 are used, while for the second message pair, the 2nd, 4th and 6th channel uses at User 1 and the 1st, 3rd and 5th channel uses at User 2 are used, as discussed in (74). In Fig. 12(11(b)), each User $i$ has $M_{i}=L_{i}/B_{i}=2$ message symbols, and the two message pairs are transmitted over 18 channel uses via alternate channel uses. Overall, through this strategy, both the users can fully utilize all channel uses.
@@
Fig. 13 shows the power distribution from RNN-based coding. We consider that the two users convey $L_{1}=L_{2}=6$ bits to one another over $N_{\text{L}}=18$ channel uses, where $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB. We consider two different coding architectures with different numbers of processing bits. First, in Fig. 13(12(a)), we consider $K_{1}=K_{2}=3$ bits, where every pair of the three bits is the input to the RNN-based coding architecture with the sigmoid function at a time to generate transmit symbols for 9 channel uses. In Fig. 13(12(b)), we consider $K_{1}=K_{2}=6$ bits, which are the input to the RNN-based coding architecture with the softmax function to generate transmit symbols for 18 channel uses.
For linear and RNN-based coding under this asymmetric SNR scenario, User 1 allocates more power at the beginning (over 6 channel uses in Fig. 12(11(a)), 9 channel uses in Fig. 13(12(a)), and 18 channel uses in Fig. 12(11(b)) and Fig. 13(12(b))), which aligns with the result of using the feedback scheme in feedback-enabled GOWCs [10, 17]. This implies that User 1 exploits the feedback information provided by User 2, in order to improve $\text{BLER}_{1}$ while sacrificing $\text{BLER}_{2}$ as shown in Fig. 9. It is worth mentioning that, as compared with the one-way feedback coding in [9, 10, 14, 17], two-way coding requires that User 2 should (i) convey its own bit vector as well as the feedback information to User 1 and (ii) satisfy the power constraint. These additional aspects are the motivation of developing novel linear and RNN-based two-way coding schemes in this work.

### v2

@@ VI-D Medium/Long Block-Lengths
(a) Sum-BER with medium/long block-lengths.
(b) Sum-BLER with medium block-length ($L_{1}=L_{2}=120$).
Figure 8: Sum-error with medium/long block-lengths for $L_{1},L_{2}$ bits where $\text{SNR}^{\text{ch}}_{1}=-1$dB and rate $r_{1}=r_{2}=1/3$. Under the high noise scenario with $\text{SNR}^{\text{ch}}_{1}=-1$dB, RNN-based coding performs well in terms of balancing the communication reliability when the channel noises become more asymmetric.
We next analyze a medium and long block-length regime, where we consider turbo coding [34] adopted in LTE standards as another baseline for medium/long block-length codes. We consider the trellis with $(4,[13,15])$, BPSK modulation, and $10$ decoding iterations. Fig. 8(8(a)) shows sum-BER curves under varying $\text{SNR}^{\text{ch}}_{2}$ where $\text{SNR}^{\text{ch}}_{1}=-1$dB and $r_{1}=r_{2}=1/3$. Since turbo coding exploits the block-length gain, a longer block of bits leads to a better sum-BER performance. However, for given $L_{1}$ and $L_{2}$, the sum-BER performances are almost constant over $\text{SNR}^{\text{ch}}_{2}$ since sum-BER is dominated by a fixed value of $\text{BER}_{1}$. On the other hand, our proposed two-way coding schemes, i.e., linear and RNN-based coding, significantly improve $\text{BER}_{1}$ as $\text{SNR}^{\text{ch}}_{2}$ increases, due to the two-way interactions between the users, which leads to substantial improvements in the sum-BER.
Fig. 8(8(b)) shows the sum-BLER performance for medium block-length with $L_{1}=L_{2}=120$, where $\text{SNR}^{\text{ch}}_{1}=-1$dB and rate $r_{1}=r_{2}=1/3$. Since the SNR condition is poor, i.e., $\text{SNR}^{\text{ch}}_{1}=-1$dB, turbo coding does not perform well, resulting in a performance degradation in terms of sum-BLER. However, the two-way coding schemes, in particular RNN-based coding, improve the error performance due to their robustness to channel noises, and ultimately balance the communication reliability, i.e., obtain much smaller sum-BLER values.
@@ VI-E Threshold for Two-Way Coding Benefits
Figure 9: Each of BLER curves obtained by RNN-based coding with the sigmoid function and $K_{1}=K_{2}=3$ bits. The users start to benefit from two-way coding when $\text{SNR}^{\text{ch}}_{2}$ is larger than a threshold, which is 6dB in (a) and 5dB in (b), respectively.
@@
Next, we investigate each error curve to analyze how communication reliability can be balanced by our two-way coding scheme under asymmetric channels. Fig. 9 shows the BLER curves obtained by our RNN-based coding scheme with the sigmoid function and $K_{1}=K_{2}=3$ bits, where BLER is measured over the $K_{1}=K_{2}=3$ bits. We consider $\text{SNR}^{\text{ch}}_{1}=1$dB in Fig. 9(9(a)) and $\text{SNR}^{\text{ch}}_{1}=-1$dB in Fig. 9(9(b)). To understand the interrelation between the two users through the proposed two-way coding, we introduce a baseline by modifying our RNN-based coding to generate two independent open-loop codes. Specifically, in Fig. 4, we provide $x_{i}[k-1]$ as an input to the encoder of User $i$ rather than $y_{i}[k-1]$ to remove the dependency between the encoders of the users, which is denoted by “RNN, open-loop” in the plots. Under the asymmetric channels with $\text{SNR}^{\text{ch}}_{1}<\text{SNR}^{\text{ch}}_{2}$, $\text{BLER}_{1}$ and $\text{BLER}_{2}$ obtained by “RNN, open-loop” is the upper and lower bound, respectively, of those obtained by our two-way coding scheme. In the figures, $\text{BLER}_{1}$ with open-loop coding yields a constant value, since $\text{BLER}_{1}$ is only determined by the fixed value of $\text{SNR}^{\text{ch}}_{1}$, while $\text{BLER}_{2}$ with open-loop coding decreases along $\text{SNR}^{\text{ch}}_{2}$. We note that the two independent open-loop codes do not balance the communication reliability when the channels are asymmetric.
@@
From the figures, we observe that our two-way coding scheme balances the communication reliability under asymmetric channels due to the interactive exchange of information between the users. $\text{BLER}_{1}$ is improved relative to that of the open-loop codes, while $\text{BLER}_{2}$ is sacrificed as compared to the open-loop coding. This is because User 2 (with higher SNR) helps User 1 (with lower SNR) through the interactive exchange of information to improve $\text{BLER}_{1}$ by implicitly providing feedback information to User 1. This behavior is in line with the insight discussed in [6] that a user with lower channel noise can act as a helper. Specifically, we can identify a threshold of $\text{SNR}^{\text{ch}}_{2}$, above which the users start to benefit from the two-way interaction in minimizing sum-errors. When $\text{SNR}^{\text{ch}}_{1}=1$dB, the threshold is $\text{SNR}^{\text{ch}}_{2}=6$dB as indicated in Fig. 9(9(a)), whereas, the threshold is $\text{SNR}^{\text{ch}}_{2}=5$dB in Fig. 9(9(b)) when $\text{SNR}^{\text{ch}}_{1}=-1$dB. Below the threshold, the BLERs of the users match the open-loop BLERs, which implies that User 2 does not provide feedback to User 1.
We also observe that $\text{BLER}_{2}$ increases slightly at the threshold and then decreases afterwards, while $\text{BLER}_{1}$ starts to decrease at the threshold. Since our objective is to minimize the “sum” of the BLERs, this behavior in the asymmetric channel indicates that User 2 begins to sacrifice its performance to improve User 1’s performance. This may lead to a slight increase in $\text{BLER}_{2}$ even with a higher SNR; for example, $\text{BLER}_{2}$ at $\text{SNR}^{\text{ch}}_{2}=6$dB is larger than at $\text{SNR}^{\text{ch}}_{2}=5$dB as shown in Fig. 9(9(a)). Overall, it is important to note that sum-BLER (the dashed black line) starts to substantially decrease at the threshold. In Fig. 9(9(a)), $\text{BLER}_{1}$ is even lower than OL-LB 1.999While we acknowledge that the lower bound, OL-LB 1, is intended for the non-cooperative case, it still provides valuable insights into the extent to which error performance can be improved through cooperation. This demonstrates that cooperation between the users enables User 1 to achieve a BLER lower than the lower bound achieved when cooperation is not allowed.
@@ VI-F The Superior Scheme over Different SNR Regions
Figure 10: The superior scheme in sum-BLER for $L_{1}=L_{2}=6$ bits under various SNR scenarios. When SNRs are high and asymmetric, linear coding performs best, while RNN-based coding excels in low SNRs.
@@
Next, we explore different coding scheme among repetitive coding, TBCC, linear coding, and RNN-based coding in terms of sum-BLER for $L_{1}=L_{2}=6$ under various SNR scenarios. The results are demonstrated in Fig. 10. In high and asymmetric SNR regions, we observe that linear coding outperforms the other schemes. For moderate and asymmetric SNRs, RNN-based coding with the sigmoid function performs the best among the schemes. In low SNR regions, RNN-based coding with the softmax function and $K_{1}=K_{2}=6$ bits yields the best performance, demonstrating its robustness to channel noises. At low/medium SNR, the noise component is relatively large, making the channel environment highly stochastic. Learning-based coding helps to extract features from this high-stochastic environment, while linear coding may not adapt well to high noise levels due to the limited degrees of freedom in coding caused by its simplified linear assumptions. At high SNR, the noise component is relatively small, making the channel environment more deterministic. Learning-based coding employing deep neural networks, such as in our RNN-based coding, may introduce unnecessary complexity in such straightforward conditions. In contrast, linear coding techniques often offer analytical solutions, leading to potentially better performance in low-randomness channel environments. Under symmetric SNR scenarios ($\text{SNR}^{\text{ch}}_{1}=\text{SNR}^{\text{ch}}_{2}=-5,-1,1,5$dB), RNN-based coding behaves like two independent open-loop coding when the channels are symmetric, as depicted in Fig. 9, which outperforms TBCC. This result suggests that our learning architecture can be also utilized to develop error correction codes. While our study focuses on two-way coding behavior, the construction of error correcting codes is a potential area for future research.
VI-G Varying Rates
@@
Figure 11: Sum-BER with various rates of $r_{2}$, where $\text{SNR}^{\text{ch}}_{1}=1$dB, $\text{SNR}^{\text{ch}}_{2}=20$dB, and $r_{1}=1/3$. RNN-based coding outperforms the counterparts in sum-BER even while supporting high rates.
@@
We so far looked into the benefit of two-way coding in terms of improving sum-error performance when rates $r_{i}$ are fixed. However, when $\text{SNR}^{\text{ch}}_{2}$ is high enough, while helping User 1 to improve sum-error, User 2 can increase its rate $r_{2}$. To investigate the relationship between the rates $r_{i}$ and sum-error, we show the sum-BER performances with varying rates of $r_{2}=K_{2}/N$ in Fig. 11, where $r_{1}=1/3$, $\text{SNR}^{\text{ch}}_{1}=1$dB, and $\text{SNR}^{\text{ch}}_{2}=20$dB. For linear coding, we consider three different modulation bits at User 1, $K_{1}=1,2,3$. Since $N=K_{1}/r_{1}$ with $r_{1}=1/3$, we have $N=3,6,9$ for $K_{1}=1,2,3$, respectively. Given $N$, we set the modulation bits at User 2 as $K_{2}=Nr_{2}$. We consider varying $r_{2}$ from $1/3$ to $4$ in the simulation. [^10] The rate $r_{i}$ in this work is defined as the number of bits per channel use. Our coding framework aligns better with coded modulation frameworks [35] because it integrates both modulation and channel coding, particularly by incorporating cooperation into coding. We consider the strategy of alternate channel uses discussed in Sec. III-H for linear coding, where two message pairs are exchanged in an alternate manner. Specifically, $2K_{1}$ and $2K_{2}$ bits are modulated to two messages at User $1$ denoted by $m_{1},m_{1}^{\prime}$ and at User $2$ denoted by $m_{2},m_{2}^{\prime}$, respectively, and the two pairs of messages, $(m_{1},m_{2})$ and $(m_{1}^{\prime},m_{2}^{\prime})$, are exchanged over $2N$ channel uses in an alternate manner. For RNN-based coding, the number of bits input to the encoders of Users 1 and 2 are $K_{1}=3$ and $K_{2}=Nr_{2}=9r_{2}$, respectively. From the figure, we observe that RNN-based coding outperforms the counterparts even under very high rates, $r_{2}\leq 3$, while linear coding outperforms the repetitive coding when the rate is quite low, $r_{2}\leq 1.5$.
@@ VI-H Power Distribution in Linear and RNN-Based Coding
Next, we investigate the power distribution of linear coding and RNN-based coding. We consider that two users convey $L_{1}=L_{2}=6$ bits to one another over $N_{\text{L}}=18$ channel uses, where $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB. We consider $K_{1}=K_{2}$. We note that the power distribution of linear coding and RNN-based coding differs based on the number of processed bits at a time, i.e., $K_{1}$ and $K_{2}$. In linear coding, $K_{i}$ bits are modulated into a message, and through the alternate channel use strategy, two message pairs are exchanged over $2N=2N_{\text{L}}/L_{i}=6K_{i}$ channel uses. In RNN-based coding, $K_{i}$ bits are input to the RNN-based coding architecture, which generates $N=N_{\text{L}}/L_{i}=3K_{i}$ transmit symbols. This means that a message pair each containing $K_{i}$ bits is exchanged over $N=3K_{i}$ channel uses. Therefore, depending on the processed bits at a time ($K_{i}$), the utilization of the channel uses will be different, which is reflected in the power distribution over the channel uses.
@@
First, we consider linear coding, where we adopt the strategy of alternate channel uses, discussed in Sec. III-H. We consider two different processing bits, $K_{1}=K_{2}=1$ in Fig. 12(12(a)) and $K_{1}=K_{2}=3$ in Fig. 12(12(b)). First, for the case of $K_{1}=K_{2}=1$ in Fig. 12(12(a)), each User $i$ has $M_{i}=L_{i}/K_{i}=6$ message symbols, and there are 6 message pairs to be exchanged. Through the alternate channel use strategy, every two message pairs are transmitted over 6 channel uses, and the pattern of power distribution during the 6 channel uses are repeated three times to exchange the 6 message pairs. Specifically, during the first 6 channel uses, to exchange the first message pair, User 1 uses the 1st, 3rd, and 5th channel uses, and User 2 uses the 2nd, 4th, and 6th channel uses, while to exchange the second message pair, User 1 uses the 2nd, 4th, and 6th channel uses, and User 2 uses the 1st, 3rd, and 5th channel uses, as discussed in (75). We observe that User 1 slightly decreases its power across the 6 channel uses, while User 2 also decreases its power across the channel uses except when transmitting its message in the 1st and 6th channel uses. For the case of $K_{1}=K_{2}=3$ in Fig. 12(12(b)), each user has $M_{i}=L_{i}/K_{i}=2$ message symbols, and the two message pairs are transmitted over 18 channel uses via alternate channel uses. Since the alternate channel use strategy is used, to exchange the first message pair, User 1 uses odd-numbered channel uses, and User 2 uses the even-numbered channel uses, while to exchange the second message pair, User 1 uses odd-numbered channel uses, and User 2 uses even-numbered channel uses. Similarly, Users 1 and 2 decrease their power across channel uses, with User 2 using the first and last channel uses to transmit its two respective messages.
@@
Next, we consider RNN-based coding, where the softmax function is used for decoding. We consider two different numbers of processing bits, $K_{1}=K_{2}=3$ bits in Fig. 13(13(a)) and $K_{1}=K_{2}=6$ bits in Fig. 13(13(b)). For the case of $K_{1}=K_{2}=3$ bits in Fig. 13(13(a)), a pair of 3 bits is input to the RNN-based coding architecture at a time to generate transmit symbols for the 9 channel uses, and this process is repeated twice to exchange a total of $L_{1}=L_{2}=6$ bits. In Fig. 13(13(a)), we observe that during the 9 channel uses, User 1 decreases its power. The behavior of User 2’s power is less interpretable due to the complexity of analyzing neural networks; however, it is clear that User 2 utilizes the 9 channel uses to provide feedback information as well as to transmit its message. For the case of $K_{1}=K_{2}=6$ bits in Fig. 13(13(b)), a pair of 6 bits is input to the RNN-based coding architecture to generate transmit symbols for 18 channel uses. Similarly, we observe that User 1 decreases its power across the channel uses. The main difference in power distribution between RNN-based coding and linear coding is the repetition pattern. As an example, we consider RNN-based coding in Fig. 13(13(a)) and linear coding in Fig. 12(12(b)), both with $K_{1}=K_{2}=3$. In RNN-based coding in Fig. 13(13(a)), the power distribution pattern repeats every $N=9$ channel uses. In contrast, in linear coding in Fig. 12(12(b)), the repeated pattern is observed in alternating channel uses over $182N=$ 18 channel uses. The difference arises because, in RNN-based coding, the two messages are transmitted successively, whereas in linear coding, the two messages are transmitted alternately over the channel uses through the alternate channel use strategy.
The key message conveyed from the power distribution results in Figs. 12 and 13 is that the helper concept in two-way channels [6] can be related to the transmitter-receiver relationship in feedback-enabled one-way channels [9, 8, 10, 15, 18] in terms of power utilization. Specifically, from the power distribution of linear and RNN-based coding under this asymmetric SNR scenario ($\text{SNR}^{\text{ch}}_{1}<\text{SNR}^{\text{ch}}_{2}$) in Figs. 12 and 13, we commonly observe that User 1 allocates more power at the beginning and decreases its power across the channel uses during the transmission of a message pair (over 6 channel uses in Fig. 12(12(a)), 9 channel uses in Fig. 13(13(a)), and 18 channel uses in Fig. 12(12(b)) and Fig. 13(13(b))). Understanding that a user with lower channel noise acts as a helper [6], we see that User 2 acts as a helper to provide feedback to User 1 to improve User 1’s error performance at the expense of sacrificing its own error performance, as demonstrated in Fig. 9. This helper concept is closely related to feedback-enabled one-way channels [9, 8, 10, 15, 18], where the receiver helps the transmitter by providing feedback to improve performance. In this context, User 1 functions similarly to a transmitter as the recipient of help in feedback-enabled one-way channels, while User 2 acts as a receiver providing feedback to User 1. It is interesting to see that the power distribution of the transmitter in feedback-enabled one-way channels, as observed in [8, 10, 18], is similar to that of User 1 in two-way channels in terms of decreasing power across the channel uses. To further understand the behavior of linear and RNN-based coding, we have included additional simulations results on scatter plots of transmit symbols across channel uses in Appendix P.